Write in standard form. It is a parabola. The vertex is (0,0) and the focus is (-6,0).
Answer by lwsshak3(11628)

Write in standard form. It is a parabola. The vertex is (0,0) and the focus is (-6,0).
This is a parabola that open leftwards.
Its standard form of equation: (y-k)^2=-4p(x-h), (h,k)=(x,y) coordinates of the vertex
For given parabola:
given vertex: (0,0)
Axis of symmetry: y=0 or x-axis
From focus:
p=6 (distance from focus to vertex on the axis of symmetry)
4p=24
Equation: y^2=-24x